please solve this for me: xy2(squared)+xyz-z2(square)
x=36, y=1/6 and z=9

Answer by ikdeep(189)

xy2(squared)+xyz-z2(square)
put the respective values of x ,y and z in the equation,,,, and we get
= (36*1/6)^2 + (36*1/6*9) - (9)^2....here ^ denotes power sign
on solving brackets we get,,,,,,,
= (6)^2 + 54 - 81
= 36 + 54 - 81
= 90 - 81
= 9 ...........
Hope this will help you
Please feel free to revert back for any further queries.